OpenStudy (anonymous):

Please Help, Jon rented a car from a company that charged a daily rental fee and a mileage charge. He rented the car for 7 days and drove 400 miles and was charged $345. His friend Amanda later rented the same car for 8 days and drove 380 miles and was charged $375. What was the daily rental charge? How much did the company charge per mile? I dont even know where to begin please help.

OpenStudy (precal):

This looks like a system of equations. What type of class are you taking?

OpenStudy (precal):

I ask because you went from a quadratic to a system.

OpenStudy (precal):

you need to use x and y Let's call x = rental days y= mileage

OpenStudy (precal):

He rented the car for 7 days and drove 400 miles and was charged $345.

OpenStudy (precal):

7x+ 400y=345 7 days is 7x and is your + 400 miles is your 400y charged 345 is =345

OpenStudy (anonymous):

Just college algebra 110

OpenStudy (precal):

Can you write the second equation we will need?

OpenStudy (precal):

Is this an online course?

OpenStudy (precal):

His friend Amanda later rented the same car for 8 days and drove 380 miles and was charged $375

OpenStudy (anonymous):

Yes and I am working on my final right now and honestly I do not remember any of this

OpenStudy (precal):

You will need this statement to create the second one.

OpenStudy (precal):

I gotta go soon so can you write your second equation.

OpenStudy (anonymous):

8x+380y=375

OpenStudy (precal):

yes

OpenStudy (precal):

8x+380y=375 7x+ 400y=345

OpenStudy (precal):

If you are given multiple choices, you can sub your x and y values to find the true statements. For example is they said (1,2) or 1 day and $2 a mile 8(1)+380(2) is 768 so that would not be the solution.

OpenStudy (precal):

You can bypass solving this using algebraic methods if you are given multiple choices. Just substitute your given choices and they do have to work for both equations. Good luck on your final. Gotta go :)

OpenStudy (anonymous):

Thank you I appreciate your time.

OpenStudy (precal):

anytime

OpenStudy (imstuck):

Those equations that precal gave you..do you know how to solve them simultaneously?

OpenStudy (precal):

He/She does not know how to solve them. In that case, the best way to solve it (if given values) is to substitute them. @IMStuck